# Benchmarking Energy Calculations Using Formal Proofs

### v1

@@ 7.   Conclusions and Outlook
In this study, we developed pairwise energy calculations in Lean and compared our results with the values provided by the NIST SRSW benchmark. Our calculations agree to machine precision with the NIST reference values. To be clear, our confidence in our system does not stem from its agreement with the NIST benchmark, rather from the theorems we have proved in Lean that certify that the functions in LeanLJ have those specified mathematical properties. For instance, the pbc function guarantees that all wrapped particles lie in the interval $[-L/2,L/2]$, and the derivation of the function computing long-range corrections is validated mathematically. We assert that LeanLJ is a more reliable benchmark than the NIST SRSW, at least for the components of the benchmark we have addressed. LeanLJ could be validated even further, by adding to the list of theorems proved about the current functions.
@@
We consider it helpful to reflect on the remaining sources of uncertainty in our code – considering what we’ve verified, what could still be wrong? First, we are trusting in the axioms of mathematics, as expressed in Lean’s core; errors here might compromise Mathlib, on which we depend. Second, our approach to polymorphism exposes us to mistakes in typeclasses. For example, we link the rounding functions for Reals and floats – a subtle error in which one rounded 0.5 down while another rounded 0.5 up would not be detected by Lean (egregious errors like linking functions for round and floor would also be technically possible, so human oversight remains necessarily at this level). Third, when we developed both polymorphic and floating-point versions of the same function for illustration purposes; we sometimes mistakenly called the floating-point version in execution instead of the polymorphic version. In this case, proofs for that function are not technically connected to the execution – since Lean doesn’t require our polymorphic code design, it doesn’t flag such mistakes (these could be eliminated by not defining Float and RealLike versions of the same functions). Fourth, we are still exposed to errors in input/output (Section 5.3), and in defining system-specific parameters, such as the force field parameters; these are mitigated by the testing, but do not prevent a user from inputting incorrect parameters for calculations outside the scope of the NIST benchmarks. A fifth source would be vulnerabilities in the broader operating system in which the code is executed. Nonetheless, traditional molecular simulation have far more possibilities for errors, such that many of these concerns are not considered in typical conversations about software correctness.
@@
In our previous work, we showed Lean’s broader utility for formalizing derivations in science as math proofs [34], digitizing key results in absorption theory, thermodynamics, and kinematics. Joseph Tooby-Smith is also developing derivations in the high-energy physics field [35, 41]. These early works showcase Lean’s rigour and versatility for building a library (or libraries) of formally-verified results in diverse areas of science, facilitating rigorous verification of scientific ideas in different disciplines.
@@
Scientific computing benchmarks are typically based on human oversight and software best practices [23]; formal verification in Lean offers an even more rigorous alternative, enabling rigorous math proofs that implemented software is correct. This shift from empirical validation to formal proof introduces a new level of confidence in molecular simulations, setting the stage for more reliable and mathematically sound scientific computing.

### v2

@@ 7.   Discussion and Outlook
In this study, we developed pairwise energy calculations in Lean and compared our results with the values provided by the NIST SRSW benchmark. Our calculations agree to machine precision with the NIST reference values. To be clear, our confidence in our system does not stem from its agreement with the NIST benchmark, rather from the theorems we have proved in Lean that certify that the functions in LeanLJ have those specified mathematical properties. For instance, the pbc function guarantees that all wrapped particles lie in the interval [L/2,L/2][-L/2,L/2], and the derivation of the function computing long-range corrections is validated mathematically. We assert that LeanLJ is a more reliable benchmark than the NIST SRSW, at least for the components of the benchmark we have addressed. LeanLJ could be validated even further by adding to the list of proved theorems about current functions.
@@
We consider it helpful to reflect on the remaining sources of uncertainty in our code – considering what we have verified, what could still be wrong? First, we are trusting in the axioms of mathematics, as expressed in Lean’s core; errors here might compromise Mathlib, on which we depend. Second, our approach to polymorphism exposes us to mistakes in our RealLike type class, as we described in Section 5.2; Lean does not check to ensure that RealLike links the correct Float- and Real-type functions. Third, when we developed both polymorphic and floating-point versions of the same function for illustration purposes; we sometimes mistakenly called the floating-point version in execution instead of the polymorphic version. In this case, proofs for that function are not technically connected to the execution – since Lean does not require our polymorphic code design, it does not flag such mistakes (these could be eliminated by not defining Float and RealLike versions of the same functions). Fourth, we are still exposed to errors in input/output (Section 5.3), and in defining system-specific parameters, such as the force field parameters; these are mitigated by the testing, but do not prevent a user from inputting incorrect parameters for calculations outside the scope of the NIST benchmarks. A fifth source would be vulnerabilities in the broader operating system in which the code is executed. Nonetheless, traditional molecular simulation have far more possibilities for errors, such that many of these concerns are not considered in typical conversations about software correctness.
@@
Nonetheless, we believe that Lean can be used (in principle) to implement molecular simulation software with all the features of the established packages used in daily practice. The user interface would not need to be more complex; the proofs and theorems would all be handled in the back-end. Furthermore, this could be integrated with software frameworks like MoSDeF [54] [33], that generate input files for multiple simulation engines. This would help address reproducibility and rigour at the “outside world” layer (Fig. 7) in which the force field parameters, molecular specifications, and configuration are defined. This would also enable rigorous evaluation of traditional software using Lean software as a benchmark; MoSDeF could create inputs for both verified Lean software as well as for LAMMPS, and deviations between the respective outputs may signal a bug in LAMMPS.
In our previous work, we showed Lean’s broader utility for formalizing derivations in science as math proofs [47], digitizing key results in absorption theory, thermodynamics, and kinematics. Joseph Tooby-Smith is also developing derivations in the high-energy physics field [48, 55]. These early works showcase Lean’s rigour and versatility for building a library (or libraries) of formally-verified results in diverse areas of science, facilitating rigorous verification of scientific ideas in different disciplines. In the long term, we envision Lean being used to formalize not just interaction potentials or particle dynamics, but the very foundations of statistical mechanics itself. The mathematics of ensemble theory, such as definitions of the microcanonical (NVE) and canonical (NVT) ensembles and proofs about their properties, could be stated precisely, and then directly linked to formally verified simulation code. For instance, we imagine proving that a molecular dynamics simulation routine satisfies the conservation laws of the NVE ensemble, in the limit of infinitesimal time steps, and verifying the detailed balance condition of Monte Carlo moves.
@@
Scientific computing benchmarks are typically based on human oversight and software best practices [36]; formal Lean verification offers an even more rigorous alternative, allowing rigorous mathematical proofs that the implemented software is correct. This shift from empirical validation to formal proof introduces a new level of confidence in molecular simulations, setting the stage for more reliable and mathematically sound scientific computing.
Conclusion
This work demonstrates how formal methods can complement molecular simulations by providing proven guarantees about the properties of the molecular simulation system. While still in its early stages, this approach opens a path toward mathematically-verified simulations.